# LLMs Struggle to Reject False Presuppositions when Misinformation Stakes are High

### v1

@@ Introduction
Speakers often have underlying assumptions, taking certain information for granted. These implicit assumptions, known as presuppositions, refer to background knowledge or shared beliefs assumed to be part of the common ground between interlocutors \@BBOPcitep\@BAP\@BBN(Stalnaker, 1973)\@BBCP. Presuppositions are introduced by specific words or syntactic structures called presupposition triggers. For instance, in the statement, “As you know, we’ve completed the wall,” the factive verb ’know’ triggers the presupposition that it is true that the wall has been completed. Presupposition triggers include a wide range of elements such as factive verbs, iteratives, or definite descriptions \@BBOPcite\@BAP\@BBNLevinson (1983)\@BBCP. They play an integral role in communication, allowing speakers to rely on shared knowledge without making it explicit. However, presuppositions are not limited to reaffirming shared beliefs; they can also introduce new information into the common ground, a phenomenon known as informative presuppositions \@BBOPcitep\@BAP\@BBN(Tonhauser, 2015)\@BBCP. Such presuppositions are particularly prominent in persuasive contexts, including media and political discourse \@BBOPcitep\@BAP\@BBN(Sbisà, 2023)\@BBCP.
@@
Intriguingly, while presuppositions enhance communication efficiency, they also pose risks when the presupposed content is false – a phenomenon known as false presuppositions or presupposition failure \@BBOPcitep\@BAP\@BBN(Yablo, 2006)\@BBCP. For example, if someone were to say, “As you know, global warming is a hoax,” this would propagate misinformation, as the presupposed claim that global warming is a deception was factually incorrect. This capacity to introduce false information into the common ground makes presuppositions a powerful yet potentially harmful linguistic tool – which is further amplified by research showing that presuppositions often prove to be more persuasive than direct assertions \@BBOPcitep\@BAP\@BBN(Thoma et al., 2023; Moldovan, 2023)\@BBCP. By presenting content as shared knowledge, presuppositions distract from critical evaluation, making them highly effective at embedding disputable or misleading information \@BBOPcitep\@BAP\@BBN(Lombardi Vallauri, 2021)\@BBCP. Behavioural experiments support this: differences in accuracy and processing times show that presuppositions cause faster and shallower processing, reducing critical attention \@BBOPcitep\@BAP\@BBN(Erickson & Mattson, 1981; Schwarz, 2015; Lombardi Vallauri, 2021)\@BBCP.
@@
Given the persuasive power and potential for misinformation inherent in presuppositions, it is important to understand how they function in communication. As large language models (LLMs) become increasingly integrated into information dissemination \@BBOPcitep\@BAP\@BBN(Kessler & Wuttig, 2024)\@BBCP, it is necessary to examine how these models handle (false) presuppositions and whether they can reject them when responding to users. In this study, we therefore examine how LLMs handle false presuppositions and explore whether specific linguistic factors systematically influence their tendency to adopt them. Understanding whether LLMs process presuppositions similarly to humans can shed light on their pragmatic competence, their ability to model nuanced human communication, and their potential role in amplifying misinformation. At the same time, LLMs offer a unique opportunity to explore theoretical questions in presupposition research. As computational models trained on vast amounts of human language data, they allow for systematic testing of how presuppositions are triggered, maintained, and resolved across different communicative contexts \@BBOPcitep\@BAP\@BBN(Contreras Kallens et al., 2023)\@BBCP. However, for LLMs to serve as a meaningful tool in presupposition research, they must demonstrate a baseline competence in processing presuppositions accurately. For this reason, we focus on cases where presuppositions introduce clear misinformation, using well-established presupposition triggers and publicly known entities (politicians) (cf. Table 1). If LLMs are to be a reliable tool for investigating presupposition theories, their responses should exhibit systematic patterns comparable to those observed in human experiments.
@@
The way presuppositions interact with their linguistic environment – such as their resistance to certain embedding contexts – offers valuable insight into the complex relationship between linguistic structures and the broader context of communication \@BBOPcitep\@BAP\@BBN(Schwarz, 2019)\@BBCP. We therefore draw on key insights from (psycho-)linguistic research on presuppositions to design our experimental conditions, grounding the analysis of LLM responses in linguistic presupposition analysis. Specifically, we investigate how factors such as presupposition triggers, embedding contexts, and scenario probabilities influence LLMs’ susceptibility to false presuppositions. Given that presuppositions can “effectively” embed disputable or misleading information, their impact is particularly significant in contexts with a high potential for misinformation. One such domain is political discourse, as it often carries deeply embedded assumptions and biases \@BBOPcitep\@BAP\@BBN(Curini & Pizzimenti, 2020)\@BBCP, making it easy for misinformation to be introduced indirectly through presuppositions Therefore, we examine LLMs’ handling of false presuppositions in a political domain. This also allows us to incorporate publicly known politicians into our prompts, providing a real-world context and control over many potential factors influencing presupposition interpretation.
@@
Research on presupposition failure in NLP has largely examined simple questions in QA systems, revealing persistent challenges in handling false presuppositions \@BBOPcitep\@BAP\@BBN(Kim et al., 2021, 2023; Daswani et al., 2024; Yu et al., 2023; Srikanth et al., 2024)\@BBCP. Existing research on LLMs in political contexts has focused on how LLMs reflect political biases, rather than their ability to recognize and handle false presuppositions \@BBOPcitep\@BAP\@BBN(Feng et al., 2023; Hartmann et al., 2023; Bang et al., 2024)\@BBCP. To our knowledge, no study has yet analyzed how LLMs handle false presuppositions in a political domain and, beyond, whether certain linguistic factors contribute systematically to LLMs adopting false presuppositions. We address this gap by examining these issues through a controlled experimental approach. Ultimately, this study aims to determine whether LLMs exhibit behavior indicative of basic pragmatic competence in presupposition processing, shedding light on their potential risks in real-world communication.
@@
We conduct an experiment to investigate whether LLMs incorrectly adopt (“accommodate”) false presuppositions or are able to reject them, focusing on the linguistic conditions that may shape their responses and on political contexts where undetected misinformation is particularly impactful. In presupposition research, accommodation refers to the process where hearers adjust their knowledge to align with the speaker’s presuppositions, even if those presuppositions are false \@BBOPcitep\@BAP\@BBN(Beaver et al., 2024)\@BBCP. For human speakers, accommodation of (false) presuppositions depends strongly on plausibility and context \@BBOPcitep\@BAP\@BBN(Degen & Tonhauser, 2021)\@BBCP. A statement like “I have to take my penguin to the vet” is less likely to be accommodated than “I have to take my dog to the vet” if the hearer is uncertain about the speaker’s pet ownership. Whether or not a presupposition is accepted depends on how the hearer assesses the context – deciding either to align with the speaker’s presuppositions, potentially accepting misinformation, or to challenge it and break the flow of conversation \@BBOPcitep\@BAP\@BBN(von Fintel, 2008; Beaver et al., 2024)\@BBCP. Since this is the first large-scale study into LLMs’ understanding of presupposition context, we designed our experimental stimuli to include presupposed misinformation in the LLM prompts that was rather implausible and blatantly false. Thus, we only tested cases where humans would be expected to directly reject the false presuppositions, rather than accommodating it.
@@
To design our experimental conditions, we draw on key insights from (psycho-)linguistic research on presuppositions. A defining characteristic of presuppositions is projection, the phenomenon where a presupposition remains intact in contexts that usually cancel entailments, such as negation, questions, or modals. For example, in both “As you don’t know, we’ve completed the wall” and “Do you know, we’ve completed the wall?”, the presupposition (the wall has been completed) remains intact. Projection from such embeddings serves as a well-established diagnostic tool for identifying presuppositions \@BBOPcitep\@BAP\@BBN(Chierchia & Mcconnell-Ginet, 1990; Bender & Lascarides, 2019)\@BBCP. Building on this, psycholinguistic research has explored factors that influence projection strength, offering valuable insights into how presuppositions behave in different contexts. For instance, several studies have shown that projection strength varies among different presupposition triggers \@BBOPcitep\@BAP\@BBN(Simons, 2001; Tonhauser et al., 2018)\@BBCP, while other research highlights the impact of embedding context \@BBOPcitep\@BAP\@BBN(Allyn Smith & Hall, 2014; Sieker & Solstad, 2022)\@BBCP, and world knowledge or probability \@BBOPcitep\@BAP\@BBN(Mahler, 2020; Degen & Tonhauser, 2021)\@BBCP. These factors directly impact how presuppositions are interpreted and processed, providing the foundation for our experimental design to systematically examine how LLMs handle false presuppositions:
@@
Embedding Context: We tested three types of embeddings: questions, negations, and modals. While prior research on false presuppositions in NLP has solely focused on simple questions, our inclusion of multiple embedding types enables a more comprehensive evaluation of presuppositions \@BBOPcitep\@BAP\@BBN(Beaver et al., 2024)\@BBCP. Each presupposition trigger was embedded in these contexts and presented as polar questions, requiring the models to either verify or reject the presupposition to provide an answer.
@@
To ensure a diverse comparison of proprietary and open instruction tuned models with strong multilingual capabilities, particularly for German, we chose OpenAI’s GPT-4-o \@BBOPcitep\@BAP\@BBN(Achiam et al., 2023)\@BBCP, MistralAI’s Mistral-7B-v03 \@BBOPcitep\@BAP\@BBN(Jiang et al., 2023)\@BBCP, and Meta’s Llama-3-8B \@BBOPcitep\@BAP\@BBN(Dubey et al., 2024)\@BBCP for evaluation. Each model was tested with the same prompt three times, yielding 3312 data points.
@@ Results
Centre-Right/Right (CDU/AfD) | 4.09 | 4.21 | 91.71 | 52.61 | 33.57 | 13.82 |
@@ Discussion
Additionally, our findings have important implications for both LLM development and real-world use. A key concern is the risk of LLMs spreading misinformation. With 35.6% of Generation Z using AI tools like ChatGPT for political information \@BBOPcitep\@BAP\@BBN(Kessler & Wuttig, 2024)\@BBCP, undetected false presuppositions could mislead users and distort public opinion. Our results suggest that instruction fine-tuning should focus on more than just enriching datasets – models need explicit components to detect the subtle ways misinformation is introduced, especially in political contexts. Improving LLMs’ sensitivity to linguistic structures like presupposition triggers and political affiliations could enhance their ability to handle misinformation more effectively. Another intriguing avenue for future research arises from the strong influence of scenario probabilities and - a little less prominently - party orientation, indicating a reliance on broader context or world knowledge.

### v2

@@ Introduction
Speakers often have underlying assumptions, taking certain information for granted. These implicit assumptions, known as presuppositions, refer to background knowledge or shared beliefs assumed to be part of the common ground between interlocutors \@BBOPcitep\@BAP\@BBN(Stalnaker1973-STAP-5)\@BBCP. Presuppositions are introduced by specific words or syntactic structures called presupposition triggers. For instance, in the statement, “As you know, we’ve completed the wall,” the factive verb ’know’ triggers the presupposition that it is true that the wall has been completed. Presupposition triggers include a wide range of elements such as factive verbs, iteratives, or definite descriptions \@BBOPcite\@BAP\@BBNLevinson_1983\@BBCP. They play an integral role in communication, allowing speakers to rely on shared knowledge without making it explicit. However, presuppositions are not limited to reaffirming shared beliefs; they can also introduce new information into the common ground, a phenomenon known as informative presuppositions \@BBOPcitep\@BAP\@BBN(Tonhauser2015)\@BBCP. Such presuppositions are particularly prominent in persuasive contexts, including media and political discourse \@BBOPcitep\@BAP\@BBN(Sbisa2023-ideology_presupposition)\@BBCP.
@@
Intriguingly, while presuppositions enhance communication efficiency, they also pose risks when the presupposed content is false – a phenomenon known as false presuppositions or presupposition failure \@BBOPcitep\@BAP\@BBN(Yablo2006-kc)\@BBCP. For example, if someone from the U.S. – let’s call him Donald Trump – were to say, “As you know, we’ve completed the wall,” this would propagate misinformation, as the presupposed claim that the wall had been completed was factually incorrect \@BBOPcitep\@BAP\@BBN(trump_claims)\@BBCP. This capacity to introduce false information into the common ground makes presuppositions a powerful yet potentially harmful linguistic tool – which is further amplified by research showing that presuppositions often prove to be more persuasive than direct assertions \@BBOPcitep\@BAP\@BBN(Thomas2023_psps_more_persuasive; Moldovan_persuasive_psps)\@BBCP. By presenting content as shared knowledge, presuppositions distract from critical evaluation, making them highly effective at embedding disputable or misleading information \@BBOPcitep\@BAP\@BBN(Lombardi-Vallauri2021-oe)\@BBCP. Behavioural experiments support this: differences in accuracy and processing times show that presuppositions cause faster and shallower processing, reducing critical attention \@BBOPcitep\@BAP\@BBN(ERICKSON1981; Schwarz2015; Lombardi-Vallauri2021-oe)\@BBCP.
@@
Given the persuasive power and potential for misinformation inherent in presuppositions, it is important to understand how they function in communication. As large language models (LLMs) become increasingly integrated into information dissemination \@BBOPcitep\@BAP\@BBN(DemokratieStudie2024)\@BBCP, it is necessary to examine how these models handle (false) presuppositions and whether they can reject them when responding to users. In this study, we therefore examine how LLMs handle false presuppositions and explore whether specific linguistic factors systematically influence their tendency to adopt them. Understanding whether LLMs process presuppositions similarly to humans can shed light on their pragmatic competence, their ability to model nuanced human communication, and their potential role in amplifying misinformation. At the same time, LLMs offer a unique opportunity to explore theoretical questions in presupposition research. As computational models trained on vast amounts of human language data, they allow for systematic testing of how presuppositions are triggered, maintained, and resolved across different communicative contexts \@BBOPcitep\@BAP\@BBN(Contreras_Kallens2023-ps)\@BBCP. However, for LLMs to serve as a meaningful tool in presupposition research, they must demonstrate a baseline competence in processing presuppositions accurately. For this reason, we focus on cases where presuppositions introduce clear misinformation, using well-established presupposition triggers and publicly known entities (politicians) (cf. Table 1). If LLMs are to be a reliable tool for investigating presupposition theories, their responses should exhibit systematic patterns comparable to those observed in human experiments.
@@
The way presuppositions interact with their linguistic environment – such as their resistance to certain embedding contexts – offers valuable insight into the complex relationship between linguistic structures and the broader context of communication \@BBOPcitep\@BAP\@BBN(Schwarz2019-xz)\@BBCP. We therefore draw on key insights from (psycho-)linguistic research on presuppositions to design our experimental conditions, grounding the analysis of LLM responses in linguistic presupposition analysis. Specifically, we investigate how factors such as presupposition triggers, embedding contexts, and scenario probabilities influence LLMs’ susceptibility to false presuppositions. Given that presuppositions can “effectively” embed disputable or misleading information, their impact is particularly significant in contexts with a high potential for misinformation. One such domain is political discourse, as it often carries deeply embedded assumptions and biases \@BBOPcitep\@BAP\@BBN(curini2020searching)\@BBCP, making it easy for misinformation to be introduced indirectly through presuppositions Therefore, we examine LLMs’ handling of false presuppositions in a political domain. This also allows us to incorporate publicly known politicians into our prompts, providing a real-world context and control over many potential factors influencing presupposition interpretation.
@@
Research on presupposition failure in NLP has largely examined simple questions in QA systems, revealing persistent challenges in handling false presuppositions \@BBOPcitep\@BAP\@BBN(kim-etal-2021-linguist; kim-etal-2023-qa; daswani2024-synqa2; yu-etal-2023-crepe; srikanth-etal-2024-pregnant)\@BBCP. Existing research on LLMs in political contexts has focused on how LLMs reflect political biases, rather than their ability to recognize and handle false presuppositions \@BBOPcitep\@BAP\@BBN(feng-etal-2023-pretraining; Hartmann2023-xz; bang-etal-2024-measuring)\@BBCP. To our knowledge, no study has yet analyzed how LLMs handle false presuppositions in a political domain and, beyond, whether certain linguistic factors contribute systematically to LLMs adopting false presuppositions. We address this gap by examining these issues through a controlled experimental approach. Ultimately, this study aims to determine whether LLMs exhibit behavior indicative of basic pragmatic competence in presupposition processing, shedding light on their potential risks in real-world communication.
@@
We conduct an experiment to investigate whether LLMs incorrectly adopt (“accommodate”) false presuppositions or are able to reject them, focusing on the linguistic conditions that may shape their responses and on political contexts where undetected misinformation is particularly impactful. In presupposition research, accommodation refers to the process where hearers adjust their knowledge to align with the speaker’s presuppositions, even if those presuppositions are false \@BBOPcitep\@BAP\@BBN(sep-presupposition)\@BBCP. For human speakers, accommodation of (false) presuppositions depends strongly on plausibility and context \@BBOPcitep\@BAP\@BBN(Degen_Tonhauser_2021_prior_beliefs)\@BBCP. A statement like “I have to take my penguin to the vet” is less likely to be accommodated than “I have to take my dog to the vet” if the hearer is uncertain about the speaker’s pet ownership. Whether or not a presupposition is accepted depends on how the hearer assesses the context – deciding either to align with the speaker’s presuppositions, potentially accepting misinformation, or to challenge it and break the flow of conversation \@BBOPcitep\@BAP\@BBN(vonFintel_2008; sep-presupposition)\@BBCP. Since this is the first large-scale study into LLMs’ understanding of presupposition context, we designed our experimental stimuli to include presupposed misinformation in the LLM prompts that was rather implausible and blatantly false. Thus, we only tested cases where humans would be expected to directly reject the false presuppositions, rather than accommodating it.
@@
To design our experimental conditions, we draw on key insights from (psycho-)linguistic research on presuppositions. A defining characteristic of presuppositions is projection, the phenomenon where a presupposition remains intact in contexts that usually cancel entailments, such as negation, questions, or modals. For example, in both “As you don’t know, we’ve completed the wall” and “Do you know, we’ve completed the wall?”, the presupposition (the wall has been completed) remains intact. Projection from such embeddings serves as a well-established diagnostic tool for identifying presuppositions \@BBOPcitep\@BAP\@BBN(Chierchia1990MeaningAG; BenderLascarides_2019)\@BBCP. Building on this, psycholinguistic research has explored factors that influence projection strength, offering valuable insights into how presuppositions behave in different contexts. For instance, several studies have shown that projection strength varies among different presupposition triggers \@BBOPcitep\@BAP\@BBN(Simons_2001; Tonhauser_etal_2018)\@BBCP, while other research highlights the impact of embedding context \@BBOPcitep\@BAP\@BBN(Allyn-Smith2014-il; SiekerSolstad_2022)\@BBCP, and world knowledge or probability \@BBOPcitep\@BAP\@BBN(Mahler_2020; Degen_Tonhauser_2021_prior_beliefs)\@BBCP. These factors directly impact how presuppositions are interpreted and processed, providing the foundation for our experimental design to systematically examine how LLMs handle false presuppositions:
@@
Embedding Context: We tested three types of embeddings: questions, negations, and modals. While prior research on false presuppositions in NLP has solely focused on simple questions, our inclusion of multiple embedding types enables a more comprehensive evaluation of presuppositions \@BBOPcitep\@BAP\@BBN(sep-presupposition)\@BBCP. Each presupposition trigger was embedded in these contexts and presented as polar questions, requiring the models to either verify or reject the presupposition to provide an answer.
@@
To ensure a diverse comparison of proprietary and open instruction tuned models with strong multilingual capabilities, particularly for German, we chose OpenAI’s GPT-4-o \@BBOPcitep\@BAP\@BBN(achiam2023gpt)\@BBCP, MistralAI’s Mistral-7B-v03 \@BBOPcitep\@BAP\@BBN(jiang2023mistral)\@BBCP, and Meta’s Llama-3-8B \@BBOPcitep\@BAP\@BBN(dubey2024LLama3herdmodels)\@BBCP for evaluation. Each model was tested with the same prompt three times, yielding 3312 data points.
@@ Results
| Centre-Right/Right (CDU/AfD) | 4.09 | 4.21 | 91.71 | 52.61 | 33.57 | 13.82 | |
@@ Discussion
Additionally, our findings have important implications for both LLM development and real-world use. A key concern is the risk of LLMs spreading misinformation. With 35.6% of Generation Z using AI tools like ChatGPT for political information \@BBOPcitep\@BAP\@BBN(DemokratieStudie2024)\@BBCP, undetected false presuppositions could mislead users and distort public opinion. Our results suggest that instruction fine-tuning should focus on more than just enriching datasets – models need explicit components to detect the subtle ways misinformation is introduced, especially in political contexts. Improving LLMs’ sensitivity to linguistic structures like presupposition triggers and political affiliations could enhance their ability to handle misinformation more effectively. Another intriguing avenue for future research arises from the strong influence of scenario probabilities and - a little less prominently - party orientation, indicating a reliance on broader context or world knowledge.